A girl is having a cupcake sale. Has a fixed cost of $10 and it costs $0.50 to make each cupcake. Find the formula for the cost function C(x), Revenue R(x) if each cupcake sells for $2, and the formula for Profit P(x).
Thank you

Answer by lwsshak3(11628)

A girl is having a cupcake sale. Has a fixed cost of $10 and it costs $0.50 to make each cupcake. Find the formula for the cost function C(x), Revenue R(x) if each cupcake sells for $2, and the formula for Profit P(x).
***
let x=number of cupcakes made
C(x)=0.5x+10
R(x)=2x
P(x)=R(x)-C(x)=2x-0.5x-10=1.5x-10
P(x)=1.5x-10